### v1

@@ Abstract
Interactions between human and objects are influenced not only by the object’s pose and shape, but also by physical attributes such as object mass and surface friction. They introduce important motion nuances that are essential for diversity and realism. Despite advancements in recent kinematics-based methods, this aspect has been overlooked. Generating nuanced human motion presents two challenges. First, it is non-trivial to learn from multi-modal human and object information derived from both the physical and non-physical attributes. Second, there exists no dataset capturing nuanced human interactions with objects of varying physical properties, hampering model development. This work addresses the gap by introducing the FORCE model, a kinematic approach for synthesizing diverse, nuanced human-object interactions by modeling physical attributes. Our key insight is that human motion is dictated by the interrelation between the force exerted by the human and the perceived resistance. Guided by a novel intuitive physics encoding, the model captures the interplay between human force and resistance. Experiments also demonstrate incorporating human force facilitates learning multi-class motion. Accompanying our model, we contribute the FORCE dataset. It features diverse, different-styled motion through interactions with varying resistances. Our code, dataset, and models will be released at https://virtualhumans.mpi-inf.mpg.de/force/.
Figure 1:  FORCE , a method and dataset for synthesizing diverse, nuanced human-object interactions by modeling the physical attributes of the interaction.
@@ 1 Introduction
Synthesizing realistic motion of human-object interaction presents a significant challenge. The complexity lies in the intricate interplay between humans and objects. While previous work has primarily addressed fundamental aspects of interactions, such as object shape and location, they overlook crucial physical attributes like weight, friction, and human force [67, 25, 92, 70]. Consequently, existing models struggle with distinguishing between carrying an empty suitcase and a full one, or determining if the interaction is feasible. These nuances are essential for diversity and realism. This work aims to fill this gap by leveraging the physical attributes to synthesize nuanced human motion in these scenarios.
@@
Physics-based methods coupled with reinforcement learning [57, 56, 19, 84] demonstrate remarkable generalization for different external forces. However, they face challenges including high complexity, because distinctive policies are usually trained with tailored reward functions for navigation and different interaction tasks. This means a hybrid approach is generally adopted. Additionally, these methods exhibit limitations in providing fine-grained control, for example, the switching between left-handed carrying to using both hands.
@@
Kinematics-based methods for synthesizing human motion are more scalable. It is imperative for real-world applications such as AR/VR and gaming, where a single model it adopted for synthesizing long-term, complex interactions. However, earlier kinematic methods either disregard the surrounding environment [46, 55, 45, 42] or focus solely on static objects [25, 70, 79, 92, 49, 91]. Closest to our goal are methods [67, 25, 30, 40] that model object shapes, but overlook physical attributes of interaction. In reality, humans adapt their motion based on perceived resistance and applied force during an object interaction [47, 20]. Take the pushing action shown in Figure 1 (left-most column). When pushing around a heavier object, the human applies a greater force, naturally adjusting posture, shifting the center of mass [73, 12], and leaning forward to push against the friction. If the resistance exceeds the applied force, the object would not move and the human would have to give up the interaction because it is not feasible. Synthesizing such nuanced motion calls for a method that generalizes well to these physical attributes of the interaction.
@@
Developing a kinematic method presents multiple challenges. Firstly, it is non-trivial to reason about multi-modal human-object information, including different actions, object poses, and crucial physical attributes. The increased complexity makes it challenging to disambiguate similar human poses, leading to motion that lacks diversity and nuanced details. Secondly, when determining the feasibility of the interaction, resistance is not the sole factor. It also depends on how the human interacts. For instance, a human can more effectively carry a heavier object with two hands than with one. It is shown that naively conditioning on the resistance generates sub-optimal results (see Section 5). Additionally, there is no dataset capturing diverse daily interactions under varying physical conditions. The lack of such data hinders model development and evaluation. Challenges present even when collecting such data, when addressing issues such as object occlusion.
@@
To address these challenges, we introduce FORCE , the first kinematic method for synthesizing human-object interactions that focuses on the nuanced details in human motion. Our method is founded on a pivotal insight: Human motion is dictated by the interrelation between the force exerted by the human and the perceived resistance. Guided by a novel intuitive physics encoding derived from these crucial attributes, our model is able to synthesize a diverse spectrum of interactions. For example, under the category of “carrying”, our model can plausibly generate motions such as carrying an object, carrying an object followed by a need to drop it, or attempting to carry but encountering failure. Additionally, we enable interactive control at run-time, the style of motion can be manipulated not only by varying the resistance of the object, but also by the desired action and contact mode (left, right, two-handed).
@@
    We introduce FORCE , the first kinematic method to synthesize human-object interaction by modeling physical attributes such as resistance and the applied human force.
@@ 2 Related Work
Interactions with static objects. Synthesizing human-object interaction is a long-standing challenge in computer vision. Its evolution spanning from the early stages of non-contextual human motion synthesis [29, 2, 58, 65, 72, 64, 17] to more recent work aimed at predicting static affordances within 3D scenes [43, 61]. This research domain primarily focuses on predicting static human entities in a given scene [93, 90, 26, 94, 86, 27], while adhering to scene constraints.
Recently, despite the substantial advancements in human-object interaction learning, most existing research focuses on human interactions with static objects. Several works [75, 74] predict human motion in pre-scanned scenes [26]. They train separate modules for predicting root trajectories initially and subsequently generating full-body poses, conditioned on both the scene and the intended path. Nevertheless, the scale and motion quality of these datasets often limit the realism of generated human motion [76, 33, 38, 49]. In efforts to enhance the quality of motion generation, recent research has concentrated on static objects interactions, such as sitting and lying on chairs or sofas [25, 91, 92, 60, 52, 95], and more recently reaching motion [37].
@@
Another avenue of exploration centers around full-body grasping synthesis [70, 79, 4]. More development effort has expanded to simultaneously synthesize human and object motions after grasping an object [22, 41]. Specialized efforts are dedicated to synthesizing dexterous hand-object interactions using object motion [71, 89, 97, 59]. Additionally, there exists a parallel work focusing on reinforcement learning (RL) [16, 9]. Despite being relevant, these methods consider humans interacting with static objects only. We address more challenging scenarios of human interacting with moving objects.
Interactions with moving objects. Investigation into synthesizing human motion for object manipulation covers both physics-based and kinematic-based approaches. In the domain of physical simulations, Merel et al. [48] presented a hierarchical RL framework that synthesizes movements related to catching and carrying boxes, utilizing egocentric observations. In a recent development, Hassan et al. [28] proposed a method based on adversarial imitation learning [53, 81, 57, 5, 56, 19, 98] for box manipulation. Furthermore, Xie et al. [84] showcased the manipulation of multiple boxes by integrating a trajectory planner with RL policy control. Although physics-based methods show promise in generalizing to the physical cues, in AR/VR or gaming applications, simplicity is valued. It is preferable to use a single model for synthesizing long-term, complex interactions rather than implementing complex methods with a family of motion policies. For this reason, our work aligns with kinematics-based approaches, aiming to provide a unified solution.
Among kinematics-based methods, the Neural State Machine (NSM) [67] stands out as a powerful method capable of modeling both static interactions, such as sitting, and dynamic interactions, including object manipulation. InterDiff  [85] supports a more diverse set of objects by reasoning about interactions in the contact space. However, it primarily considers short-term interactions and assumes stable contact between human and objects. More recently, OMOMO [40] conditions on the object’s trajectory to predict full-body object manipulation, but it neglects the physical attributes of the interaction, such as the mass or friction. Our work leverages these overlooked physical attributes and synthesizes diverse human-object interactions with nuanced details.
@@
Intuitive physics guided learning. The contemporary studies of intuitive physics [47, 77, 6] has experienced a resurgence. This is attributed to recent attempts to incorporate human-level intuitive physics capabilities into deep learning frameworks. Some work infers the intrinsic physical properties of the object such as mass or spring [20, 96], while others predict the underlying object dynamics such as object trajectory or the forces [51, 50, 18, 78, 39, 36, 23], which we found to be more relevant. Some of the investigations in this field forecast object trajectories or forces using Newtonian physics principles from images [51, 50], while others predict physical stability in moving object scenarios [78, 39, 36, 23]. Applications of intuitive physics have recently expanded to 3D computer vision. These include recovering object trajectories by modeling gravity [18], and guiding 3D human pose estimation with center of mass and center of pressure.
@@
We draw our key insight from the intuitive physics studies, and model intricate interplay between dynamic human forces and object’s resistance.
3 Dataset
Our goal is to build a model of human-object interactions that is aware of the intuitive physics, yet existing datasets fall short in accurately capturing human interactions with moving objects [67, 25, 92, 8, 34]. A recent dataset offers high-quality dynamic interaction data, but it does not capture different styled human motion under varying physical conditions [40] (see Table 1). To address this gap, we introduce the FORCE dataset, which comprises $450$ motion sequences of human pushing, pulling and carrying objects. It includes accurately tracked human and objects. The dataset’s main characteristic is its diverse motion nuances through interactions with varying resistances. We capture interactions with 8 objects at 3-6 levels of resistance. We hold out 30 sequences of data covering a range of interactions and scenarios for evaluation.
Table 1: Comparison with existing human-object interaction datasets with objects. The criteria include (i) high MoCap quality (ii) capture of moving objects (iii) RGB-D (iv) nuanced interaction with varying levels of resistance.
Dataset High Quality Moving Obj. RGB-D Resistance
[…]
COUCH [92]
[…]
OMOMO [40]
Ours
Human tracking. To overcome the susceptibility to noise and object occlusion, we improve upon existing depth-only camera-based capture systems [8, 34, 26]. We leverage a customized tracker, integrating 17 human-mounted IMU [54] sensors with the cameras. This enhancement improves the accuracy of the captured data. After initially fitting the SMPL[44] to the captured point clouds, we synchronize the tracking results with motion sequences recorded by IMU sensors. Subsequently, we optimize the body poses from the IMU tracking to align with the Kinect-fitted result. Further technical details about our human tracking methodology are available in the supplementary materials.
Objects tracking. We initiate by pre-scanning eight objects (carry: bin, suitcase, stool, basket, container, backpack; push/pull: chair, box) to obtain their template meshes. Subsequently, we fit them to annotated keypoints on the captured images using camera projection. The fitting is refined by running ICP to fit to the segmented object point cloud [13]. We consider our objects to be rigid.
Diverse interactions with varying resistance. After building our customized tracker, we systematically capture interactions with objects placed in varying positions relative to the human. To ensure the authenticity of the motion, the subject is only offered high-level guidance instead of specific instructions. Figure 2 showcases a selection of interactions. The capture is designed to cover diverse factors of variation. Firstly, we vary the resistance of the object in each interaction instance with removable weights. The same motion is performed for each object. Secondly, the dataset captures different contact modes including one-handed and two-handed interactions.
Note, our data collection is performed on a uniform surface, and we determine the frictional coefficient by measuring force meter readings during the uniform pulling of a 10kg object. For consistency, we approximate gravity ($g$) and the frictional coefficient ($\eta$) as 10N/kg and 1, respectively.
Object shape augmentation. To generalize to unseen shapes, each motion sequence is augmented with 10 differently shaped objects from one of the table, chair, bag, and shelf categories of ShapeNet [10] (see the last column of Figure 2). First, we detect the ground truth hand contact positions on the object surface based on a distance threshold of 0.05cm. Next, we transfer contacts by aligning the new object with the source via centering and scaling. The ground truth contacts are projected onto the nearest surface of the new object. At every frame, following object sampling and rescaling, the human poses are recomputed to satisfy the contacts using inverse kinematics [67]. We assume the resistance remains unchanged after augmentation to preserve the context of human motion.
How does this dataset benefit the community?
  • Developing physics-based synthesis methods [28, 81, 53].
  • Synthesizing non-physics-informed object interactions [67, 25, 40, 92, 85].
  • Human-object reconstruction [82, 83, 59, 18].
  • Benchmarking pose estimation for occlusion scenarios [14, 63, 15].
  • Tracking human-object interaction with RGB-D [8, 32].
@@ 4.2 MNet: Physics-aware Motion Prediction
At current frame $i$, given the poses of the human $\mathcal{H}_{i}$ and the object $\mathbf{O}_{i}$, the object geometry $\mathbf{O}_{i}$, MNet leverages the interrelation between the applied human force and the perceived resistance to auto-regressively synthesize the future human motion ~$\tilde{\mathcal{H}}_{i+1}$. At the core of MNet, is the intuitive physics encoding $\mathcal{F}_{i}$, which encapsulates the crucial physics attributes of the interaction.
@@
Geometry $\mathbf{G}_{i}$: The object geometry $\mathbf{G}_{i}$ is encoded by voxelizing object shape (in an $8\times 8\times 8$ dimensional grid). Each voxel stores its occupancy ($\mathbb{R}$) and the relative vector between the human root joint and the voxel ($\mathbb{R}^{3}$). This allows us to reason about the shape of the object and the distance between the person and different parts of the object. We vectorize this grid to obtain our geometry encoding $\mathbf{G}_{i}\in\mathbb{R}^{2048}$.
Phase $\mathbf{\Phi}_{i}$: We introduce a learned 4-dimensional vector, phase $\mathbf{\Phi}_{i}\in\mathbb{R}^{4}$, encoding the human joint trajectories following [66]. It encapsulates the spatial-temporal context of the motion.
Training. The MNet adopts a mixture-of-experts framework [31, 25, 67, 68, 92, 91] and is trained in a supervised manner by directly minimizing the MSE loss on the following outputs.
@@
Derivation of Human Force. For training, the human force is not directly observable. However, they can be estimated from the observable object acceleration. a) When the object moves, we leverage Newton’s laws to compute the human force. We assume the direction of force acts opposite to the direction of the object acceleration. For example, for pushing and pulling, the force is derived as $m\cdot g\cdot\eta\cdot\frac{\boldsymbol{a}}{\|\boldsymbol{a}\|}+m\cdot\boldsymbol{a}$, where $m$, $g$, $\eta$, $\boldsymbol{a}$ denote the mass, gravity, frictional coefficient and object acceleration. b) When the object is in contact while remaining static, we derive the force by linearly interpolating between the zero, and the derived force from the nearest frame at which the object moves. Furthermore, to eliminate the effect of noise of object acceleration, we apply a Butterworth filter of kernel size 4, while preserving its range.

### v2

@@ Abstract
Interactions between human and objects are influenced not only by the object’s pose and shape, but also by physical attributes such as object mass and surface friction. They introduce important motion nuances that are essential for diversity and realism. Despite advancements in recent human-object interaction methods, this aspect has been overlooked. Generating nuanced human motion presents two challenges. First, it is non-trivial to learn from multi-modal human and object information derived from both the physical and non-physical attributes. Second, there exists no dataset capturing nuanced human interactions with objects of varying physical properties, hampering model development. This work addresses the gap by introducing the FORCE model, an approach for synthesizing diverse, nuanced human-object interactions by modeling physical attributes. Our key insight is that human motion is dictated by the interrelation between the force exerted by the human and the perceived resistance. Guided by a novel intuitive physics encoding, the model captures the interplay between human force and resistance. Experiments also demonstrate incorporating human force facilitates learning multi-class motion. Accompanying our model, we contribute a dataset, which features diverse, different-styled motion through interactions with varying resistances. Our code, dataset, and models will be released at https://virtualhumans.mpi-inf.mpg.de/force/.
@@ 1 Introduction
Synthesizing realistic motion of human-object interactions (HOI) presents a significant challenge. The complexity lies in the intricate interplay between humans and objects. While previous work has primarily addressed fundamental aspects of interactions, such as object shape and location, they overlook crucial physical attributes like weight, friction, and human force [98, 33, 137, 101]. Consequently, existing models struggle with distinguishing between carrying an empty suitcase and a full one, or determining if the interaction is feasible. These nuances are essential for diversity and realism. This work aims to fill this gap by leveraging the physical attributes to synthesize nuanced human motion in these scenarios.
@@
As a parallel paradigm, physics-based methods using reinforcement learning coupled with imitation learning [86, 85, 26, 119] demonstrate remarkable generalization for different scenarios. However, they face challenges in providing direct fine-grained control, for example, the switching between left-handed carrying to using both hands. Additionally, these methods are computationally demanding, because distinctive policies need to be trained with tailored reward functions for navigation and interaction tasks.
@@
Kinematic methods for human motion synthesis are known for their flexibility for fine-grained spatial or temporal motion control. They are scalable to train and easy to deploy. However, earlier approaches either disregard the surrounding environment [73, 83, 72, 63] or focus solely on static objects [33, 101, 115, 137, 77, 136]. Closest to our goal are methods [98, 33, 39, 59] that model object shapes, but overlook physical attributes of interaction. In reality, humans adapt their motion based on perceived resistance and applied force during an object interaction [74, 27]. Take the pushing action shown in Figure LABEL:fig:teaser (left-most column). When pushing around a heavier object, the human applies a greater force, naturally adjusting posture, shifting the center of mass [105, 14], and leaning forward to push against the friction. If the resistance exceeds the applied force, the object would not move and the human would have to give up the interaction because it is not feasible. Synthesizing such nuanced motion calls for a method that generalizes well to these physical attributes of the interaction.
@@
Developing such a method presents multiple challenges. Firstly, it is non-trivial to reason about multi-modal human-object information, including different actions, object poses, and crucial physical attributes. The increased complexity makes it challenging to disambiguate similar human poses, leading to motion that lacks diversity and nuanced details. Secondly, when determining the feasibility of the interaction, resistance is not the sole factor. It also depends on how the human interacts. For instance, a human can more effectively carry a heavier object with two hands than with one. It is shown that naively conditioning on the resistance generates sub-optimal results (see Section 5). Additionally, there is no dataset capturing diverse daily interactions under varying physical conditions. The lack of such data hinders model development and evaluation. Challenges present even when collecting such data, when addressing issues such as object occlusion.
@@
To address these challenges, we introduce FORCE , a method for synthesizing human-object interactions that focuses on the nuanced details in human motion. Our method is founded on a pivotal insight: Human motion is dictated by the interrelation between the force exerted by the human and the perceived resistance. Guided by a novel intuitive physics encoding derived from these crucial attributes, our model is able to synthesize a diverse spectrum of interactions. For example, under the category of “carrying”, our model can plausibly generate motions such as carrying an object, carrying an object followed by a need to drop it, or attempting to carry but encountering failure. Additionally, we enable interactive control at run-time, the style of motion can be manipulated not only by varying the resistance of the object, but also by the desired action and contact mode (left, right, two-handed).
@@
    We introduce FORCE , the first kinematic method to synthesize human-object interaction by modeling physical attributes such as resistance and the applied human force. It achieves state-of-the-art results quantitatively and qualitatively.
@@ 2 Related Work
Human Motion Synthesis. Human motion synthesis has been a long-standing challenge in computer vision, evolving from early non-contextual approaches [37, 2, 87, 95, 140, 15, 51, 55, 5], to more sophisticated methods. Recent advancements include text-to-motion [104, 94, 133, 134, 6, 16, 22, 52, 38, 65, 70, 71, 42], motion-to-text [139, 46], motion synthesis with spatial [49, 118, 106, 25] or temporal control [89], audio to motion [75, 18]. Other directions explore human-to-human interaction [30, 68] and learning 3D human motions from images and videos [122, 125, 126].
In the context of human-object interactions, research has progressed from predicting static affordances in 3D scenes [64, 91, 138, 135, 34, 142, 123, 35, 60] to dynamic human motions. However, most existing work focuses on human interactions with static scenes [111, 41, 54, 77, 34, 109, 108, 20, 57], while some studies have attempted to enhance motion generation quality by concentrating on static object interactions, such as sitting and lying on furniture [33, 136, 137, 90, 81, 143, 127, 53].
@@
A parallel paradigm of research explores reinforcement learning (RL) for object manipulation [19, 10, 11], carrying objects [36, 107, 76, 119], and performing locomotion [66, 92]. RL has also been applied to interactions with static objects [81, 117] sports movements [110] and text to motion [21]. Another avenue of exploration centers around full-body grasping synthesis [101, 115, 4, 103, 61] and simultaneous synthesis of human and object motions [29, 62, 67, 96]. Specialized efforts are dedicated to synthesizing dexterous hand-object interactions using object motion [102, 131, 145, 88].
While these approaches have made significant strides, they primarily deal with static object interactions or specific types of movements. Our work addresses more challenging scenarios involving humans interacting and moving objects.
Interactions with moving objects. Recent datasets have stimulated development in human-object interaction synthesis, capturing single-person [47, 141, 43, 9, 59, 48, 50] and multi-person [129, 132] with dynamic objects. Among the current methods for synthesizing dynamic human-object interactions, the Neural State Machine [98] can model both static and dynamic interactions. Diffusion models [24, 120, 84, 114, 121, 124] are applied to synthesize short-term human-object interactions with a fixed set of objects [9], often assuming stable contact between humans and objects. More recent work [59, 58] conditions on object trajectories or waypoints to predict full-body object manipulation.
However, these previous datasets and methods overlook how human motion is dictated by the physical attributes of the interaction, such as mass or friction. Our work addresses this gap by capturing a dataset that focuses on interactions with objects of varying resistance, including pushing, pulling, and carrying. We emphasize the nuanced details and motion diversity resulting from these properties.
@@
Intuitive physics guided learning. The study of intuitive physic [74, 112, 7] has experienced a resurgence with recent attempts to incorporate human-level intuitive physics capabilities into deep learning frameworks. Some work focuses on inferring intrinsic physical properties of objects such as mass or spring [27, 144] while others predict underlying object dynamics such as the trajectory [79, 78], forces [23, 113, 56, 45, 31], or physical stability [113, 56, 45, 31].
Applications of intuitive physics have expanded to 3D computer vision, including recovering object trajectories by modeling gravity  [23], and guiding 3D human pose estimation with center of mass and center of pressure [105].
@@
Our work draws key insights from intuitive physics studies, modeling the intricate interplay between dynamic human forces and object resistance, by explicitly incorporating these physical attributes to generate more realistic and diverse human-object interactions.
@@ 3.2 MNet: Physics-aware Motion Prediction
At current frame $i$, given the poses of the human $\mathcal{H}_{i}$ and the object $\mathbf{O}_{i}$, the object geometry $\mathbf{G}_{i}$, MNet leverages the interrelation between the applied human force and the perceived resistance to auto-regressively synthesize the future human motion ~$\tilde{\mathcal{H}}_{i+1}$. At the core of MNet, is the intuitive physics encoding $\mathcal{F}_{i}$, which encapsulates the crucial physics attributes of the interaction.
@@
Geometry $\mathbf{G}_{i}$: The object geometry $\mathbf{G}_{i}$ is encoded by voxelizing object shape (in an $8\times 8\times 8$ dimensional grid). Each voxel stores its occupancy ($\mathbb{R}$) and the relative vector between the human root joint and the voxel ($\mathbb{R}^{3}$). This enables reasoning about the shape of the object. We vectorize this grid to obtain our geometry encoding $\mathbf{G}_{i}\in\mathbb{R}^{2048}$.
Phase $\mathbf{\Phi}_{i}$: We introduce a learned variable, phase $\mathbf{\Phi}_{i}\in\mathbb{R}^{4}$, encoding the human joint trajectories following [97]. It encapsulates the spatial-temporal context of the motion.
Training. The MNet adopts a mixture-of-experts architecture [40, 33, 98, 99, 137, 136]. It is trained with direct supervision by minimizing the MSE loss on the outputs:
@@
Derivation of Human Force. For training, the human force is not directly observable. However, it can be estimated from the observable object acceleration. a) When the object moves, we leverage Newton’s laws to compute the human force. We assume the direction of force acts opposite to the direction of the object acceleration. For example, for pushing and pulling, the force is derived as $m\cdot g\cdot\eta\cdot\frac{\boldsymbol{a}}{\|\boldsymbol{a}\|}+m\cdot\boldsymbol{a}$, where $m$, $g$, $\eta$, $\boldsymbol{a}$ denote the mass, gravity, frictional coefficient and object acceleration. b) When the object is in contact while remaining static, we derive the force by linearly interpolating between the zero, and the derived force from the nearest frame at which the object moves. c) When the resistance is too great for a successful interaction, we leverage the maximal observed human force for each action class as the control signal. Note, to eliminate the effect of noise of object acceleration, we apply a Butterworth filter of kernel size 4.
@@ 3.3 Resistance-conditioned Contact Prediction
4 Dataset
We introduce a dataset that comprises $450$ motion sequences of human pushing, pulling and carrying. Its main characteristic is the diverse motion nuances through interactions with varying resistances. It includes 8 objects, captured with 3-6 levels of resistance (from 5kg to 35kg) with removable weights. The testing set consists of 30 sequences covering a range of interactions and scenarios.
Human and object tracking. We leverage a customized tracker, integrating 17 human-mounted IMU [82] sensors with the cameras. This enhancement improves the accuracy of the captured data. After initially fitting the SMPL[69] to the captured point clouds, we synchronize the tracking results with motion sequences recorded by IMU sensors. Subsequently, we optimize the body poses from the IMU tracking to align with the Kinect-fitted result.
For objects, we initiate the tracking by pre-scanning eight objects (carry: bin, suitcase, stool, basket, container, backpack; push/pull: chair, box) to obtain their template meshes. Subsequently, we fit them to annotated keypoints on the captured images using camera projection. The fitting is refined by running ICP to fit to the segmented object point cloud [17]. We consider our objects to be rigid.
Further technical details about our tracking methodology are available in the supplementary materials.
Diverse interactions with varying resistance. The capture covers diverse motion variation. Firstly, we vary the resistance of the object in each interaction instance. The same motion is performed for each object of the same action category. Secondly, the dataset captures different contact modes including one-handed and two-handed interactions. We systematically place them in varying positions relative to the human. To ensure the authenticity of the motion, only high-level guidance is provided instead of specific instructions. Figure 2 showcases a selection of interactions.
Note, that our data collection is performed on a uniform surface, and we determine the frictional coefficient by measuring force meter readings during the uniform pulling of a 10kg object. For consistency, we approximate gravity ($g$) and the frictional coefficient ($\eta$) as 10N/kg and 1, respectively.
Figure 3: Comparison with baselines. Our intuitive physics encoding $\mathcal{F}$ disambiguates different multi-class motions and enables the synthesis of the desired interaction (carrying in a, pulling in b). c) and d) show the coupled encoding of human force and resistance allows the synthesis of visually plausible interactions. In c), when the resistance is greater than the human force, FORCE synthesizes a more plausible “infeasible” interaction. In d), with low resistance, the human pose maintains an upright position.
Object shape augmentation. To generalize to unseen shapes, each motion sequence is augmented with 10 differently shaped objects from one of the table, chair, bag, and shelf categories of ShapeNet [12] (see the last column of Figure 2). First, we detect the ground truth hand contact positions on the object surface based on a distance threshold of 0.05cm. Next, we transfer contacts by aligning the new object with the source via centering and scaling. The ground truth contacts are projected onto the nearest surface of the new object. At every frame, following object sampling and rescaling, the human poses are recomputed to satisfy the contacts using a full-body inverse kinematics [98]. We assume the resistance remains unchanged after augmentation to preserve the context of human motion.